Given
f(x) = 2x2 + 1
and
g(x) = 8x − 9,
find the following.
(a)
f(g(2))



(b)
f(g(x))


(c)
g(f(x))


(d)
(g of g)


(e)
(f 0f f)(-2)

Answer by stanbon(75887)

Given
f(x) = 2x^2 + 1
and
g(x) = 8x − 9,
find the following::
(a)
f(g(2)) = f(8*2-9) = f(7) = 2*7^2+1 = 2*49+1 = 99
------------------------
(b)
f(g(x)) = f(8x-9) = 2(8x-9)^2 + 1 = etc.
-----------------
(c)
g(f(x)) = g(2x^2+1) = 8(2x^2+1)-9 = 16x^2 -1
---------------------
(d)
(g of g) = g(g(x) = g(8x-9) = 8(8x-9)-9 = 64x-81
-------------------------
(e)
(fof)(-2) = f(f(-2)) = f(2*(-2)^2+1) = f(9) = 2*9^2+1 = 163
----------
Cheers,
Stan H.
------------